OpenStudy (anonymous):

A compound contains only carbon and hydrogen. A 0.857 g sample is burned in oxygen to give 1.45 L of carbon dioxide, measured dry at 98.0 kPa and 20.0 degrees C. What is the compound's empirical formula?

OpenStudy (anonymous):

This is the formula to calculate the moles of carbon dioxide: \[n=\frac{PV}{RT}\] These are the values of the variables: \[R=8.31\frac{kPaL}{molK}\] \[T=20C=293.15K\] \[P=98.0kPa\] \[V=1.45L\] Plug them into the formula: \[n=\frac{98.0*1.45}{8.31*293.15}=0.0583 moles CO_2\] Calculating Carbon moles: \[0.0583molesCO_2*\frac{1moleC}{1molesCO_2}=0.0583molesC\] Now we can use the molecular weight to determine how much of the weight represents carbon: \[0.0583molesC\frac{12.0107gC}{1moleC}=0.701gC\]

OpenStudy (anonymous):

Calculating the amount of hydrogen: \[0.857g-0.701g=0.156g\] \[0.156gH\frac{1.0079gH}{1moleH}=0.158molesH\] Divide by lowest value to determine ratios: \[\frac{0.158}{0.0583}\approx2.7\] Rounding up we can determine that we have: \[CH_3\]

OpenStudy (anonymous):

Woops. I made a booboo on the second set of calculations. Should be: \[0.156gH\frac{1moleH}{1.0079gH}=0.155molesH\] Note the grams have to cancel. \[\frac{0.155}{0.0583}=2.65\approx3\] The answer still ends up being \[CH_3\] This is based upon feedback I received from my Chem teacher during summer. I was advised that you aim for the smallest set if variables as large molecules usually don't exist in nature.